specific volume of cylindrical virus particle 6.02/100cc/gm whose radiud and length are7 angstrom and 10angstrom respectively find mol.wt. of virus
 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Expression for specific volume is incorrect or unclear,

However it is reciprocal of density

Volume times density will give mass of the cylindrical virus, hope this helps

Regards
Arun
4 years ago
Volume of one virus = voume of cylinder = π r²h 
= 22/7 × 7 × 7 × 10 A°³
=22 × 7 × 10 A°³
=1540 × 10^-30 m³ 
=1.54 × 10^-27 m³ 
so, volume of one mole of virus = 6.023 × 10²³ × volume of one virus 
= 6.023 × 10²³ × 1.54 × 10^-27 m³ 
= 9.27542 × 10^-4 m³
= 9.27542 × 10² cm³ 
now, molecular weight = volume of 1mole/ specific volume 
= 9.27542 ×10² / 6.02 × 10^-2 g/mol 
= 1.54 × 10⁴ g/mol